OpenStudy (anonymous):

elaine opened her science book to the first two pages of chapter 7. The sum of the facing pages was 437. What were the page numbers of the first two pages of chapter 7?

OpenStudy (anonymous):

this is simple all you need to do is divide by 2 and then add .5 to one and subtract to the other so 437 /2 = 218.5 so the pages are 218 and 219

OpenStudy (blacksteel):

We know that the page numbers are going to be 1 apart from each other. Let's let the number of the first page be x. Then the number of the second page is x+1. Then x + x + 1 = 2x + 1 = 437 2x = 436 x = 218 So the first page number is x = 218, and the second page number is x + 1 = 219.

OpenStudy (anonymous):

um .....goodjob blacksteel...